OpenStudy (anonymous):

Identify the 42nd term of an arithmetic sequence where a_1 = −12 and a_27 = 66.

OpenStudy (anonymous):

A. 70 B. 72 C. 111 D. 114

hartnn (hartnn):

do you know the n'th general term formula for arithmetic sequence ?

OpenStudy (anonymous):

No but I think the answer is C. 111

hartnn (hartnn):

did you calculate it ?

OpenStudy (anonymous):

The best I know how... I'm not sure if I have done it right though a(n) = a(1) + (n - 1)d a(1) = -12 a(n) = -12 + (n - 1)d a(27) = 66 66 = -12 + (27 - 1)d 66 = -12 + 26d 26d = 66 + 12 26d = 78 d = 78/26 d = 3 a(n) = -12 + 3(n - 1) a(42) = -12 + 3(42 - 1) = -12 + 3(41) = -12 + 123 = 111

hartnn (hartnn):

oh wow11 thats correct :)

OpenStudy (anonymous):

Awesome! thank you :)

hartnn (hartnn):

welcome ^_^